# Model Alignment Search

### v5

@@ A.1 Model Details
All artificial neural network models were implemented and trained using PyTorch [32] on a single Nvidia Titan X GPU. Unless otherwise stated, all models used an embedding and hidden state size of 128 dimensions. To make the token predictions, each model used a two layer multi-layer perceptron (MLP) with GELU nonlinearities, with a hidden layer size of 4 times the hidden state dimensionality with 50% dropout on the hidden layer. The GRU and LSTM model variants each consisted of a single recurrent cell followed by the output MLP. Unless otherwise stated, the transformer architecture consisted of two layers using Rotary positional encodings [37]. Each model variant used the same learning rate scheduler, which consisted of the original transformer [40] scheduling of warmup followed by decay. We used 100 warmup steps, a maximum learning rate of 0.001 , a minimum of 1e-7, and a decay rate of 0.5. We used a batch size of 128, which caused each epoch to consist of 8 gradient update steps.
A.2 MAS (and Associated Variants) Training Details
@@
All MAS trainings were implemented and trained using PyTorch [32] on single Nvidia Titan X GPUs. For each rotation matrix training, we use 10000 intervention samples and 1000 samples for validation and testing. We uniformly sampled corresponding indices upon which to perform interventions, excluding the B, T, and E tokens in the numeric equivalence tasks from possible intervention sample indices. In the Arithmetic task, we used the comma token for Rem Ops and Cumu Val interventions. When intervening upon a state in the demo phase in the numeric equivalence tasks, we uniformly sample a number of steps to continue the demo phase that will keep the object quantity below 20. We orthongonalize the matrices, $Q_{i}$, using PyTorch’s orthogonal parametrization with default settings. PyTorch creates the orthogonal matrix as the exponential of a skew symmetric matrix. We train the rotation matrices for 1000 epochs, with a batch size of 512 used for each model index pairing. We only perform experiments considering two models. Each gradient step uses the average gradient over batches of all 4 $i,j$ pairings. We select the checkpoint with the best validation performance for analysis. We use a learning rate of 0.001 and an Adam optimizer.
A.3 RSA Details
@@
We performed RSA on a subsample of a dataset of 15 sampled sequences for each object quantity ranging from 1-20 on the task that each model was trained on for each model. We first ran the models on their respective datasets to collect the latent representations. We sampled 1000 of these latent vectors as the sample representations in a matrix $M_{k}\in R^{N\times d_{k}}$ where $k$ refers to the model index, $N$ is the number of latent vectors ($N=1000$ in our analyses), $d_{k}$ is the dimensionality of a single latent vector for model $k$. We then calculated the sample cosine distance matrices (1-cosine similarity) for each model resulting in matrices $C_{k}\in R^{N\times N}$. Lastly we calculated the Spearman’s Rank Correlation Coefficient between the lower triangles of the matrices $C_{1}$ and $C_{2}$ as the RSA value using python’s SciPy package [46, 25, 41]. We resampled the 1000 vectors 10 times and recalculated the RSA score 10 times and report the average over these scores.
A.4 CKA Details
@@
[24]. We resampled the 1000 vectors 10 times and recalculated the CKA score 10 times and report the average over these scores.
A.5 DeepSeek
To avoid fine-tuning the DeepSeek model, we prompted the model to perform a numeric equivalence task similar to the Multi-Object task. The prompt consisted of: """ User: For each trial, I am going to say "nut" some number of times and then I want you to respond by saying "seed" the exact same number of times that I said "nut" in that trial. Please say "stop" at the end of your response to indicate that you are finished. Do not say or think anything else. Example: User: nut : seed stop Example: User: nut nut : seed seed stop Begin Trial: User: """ Following the prompt, we provided some number of "nut" tokens and considered a trial correct when DeepSeek responded with the corred number of "seed" tokens followed by the "stop" token. DeepSeek was only successful on counts up to 12, so the MAS analysis only included values from 1-12.
@@ A.6 Arithmetic
[…]
@@ A.7 Symbolic Variables
A.8 Gram-Schmidt Intervention Proof
We set out to prove the following:
$\mathcal{R}_{i}^{-1}((1-D_{i})\mathcal{R}_{i}h^{(i)}_{trg,t}+D_{j}\mathcal{R}_{j}h^{(j)}_{src,u})=h^{(i)}_{trg,t}+Q_{i}(-Q_{i}^{\top}h^{(i)}_{trg,t}+Q_{j}^{\top}h^{(j)}_{src,u})$ (20)
Where $i$ and $j$ indicate the model indices, ${src}$ and ${trg}$ indicate which is the source of information and which is the receiver of information, $t$ and $u$ are integers referring to the step in the target and source sequences from which the tokens were obtained, $\mathcal{R}_{i}\in R^{d\times d}$ is an orthonormal matrix, $D_{i}\in R^{d\times d}$ is a diagonal binary matrix with $d_{var}$ nonzero elements, each $h\in R^{d}$ is a latent state vector, and $Q\in R^{d\times d_{var}}$ is a matrix composed of $d_{var}$ orthogonal column vectors of unit length. We also assume that $\begin{bmatrix}Q_{k}^{\top}\\ 0\end{bmatrix}=D_{k}\mathcal{R}_{k}$ where $D_{k}=\begin{bmatrix}I&0\\ 0&0\end{bmatrix}$ and $I\in R^{d_{var}\times d_{var}}$ is an identity matrix. The intervened vector $v_{trg,t}^{(i)}$ for a given intervention is calculated as follows:
$v^{(i)}_{trg,t}$ $=$ $\mathcal{R}_{i}^{-1}((1-D_{i})\mathcal{R}_{i}h^{(i)}_{trg,t}+D_{j}\mathcal{R}_{j}h^{(j)}_{src,u})$ (21)
$v^{(i)}_{trg,t}$ $=$ $\mathcal{R}_{i}^{\top}(1-D_{i})\mathcal{R}_{i}h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}D_{j}\mathcal{R}_{j}h^{(j)}_{src,u}$ (22)
$v^{(i)}_{trg,t}$ $=$ $\mathcal{R}_{i}^{\top}\mathcal{R}_{i}h^{(i)}_{trg,t}-\mathcal{R}_{i}^{\top}D_{i}\mathcal{R}_{i}h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}D_{j}\mathcal{R}_{j}h^{(j)}_{src,u}$ (23)
$v^{(i)}_{trg,t}$ $=$ $h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}(-D_{i}\mathcal{R}_{i}h^{(i)}_{trg,t}+D_{j}\mathcal{R}_{j}h^{(j)}_{src,u})$ (24)
$v^{(i)}_{trg,t}$ $=$ $h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}(-\begin{bmatrix}Q_{i}^{\top}\\ 0\end{bmatrix}h^{(i)}_{trg,t}+\begin{bmatrix}Q_{j}^{\top}\\ 0\end{bmatrix}h^{(j)}_{src,u})$ (25)
$v^{(i)}_{trg,t}$ $=$ $h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}(-\begin{bmatrix}I\\ 0\end{bmatrix}Q_{i}^{\top}h^{(i)}_{trg,t}+\begin{bmatrix}I\\ 0\end{bmatrix}Q_{j}^{\top}h^{(j)}_{src,u})$ (26)
$v^{(i)}_{trg,t}$ $=$ $h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}\begin{bmatrix}I\\ 0\end{bmatrix}(-Q_{i}^{\top}h^{(i)}_{trg,t}+Q_{j}^{\top}h^{(j)}_{src,u})$ (27)
Here we note that if we set $S_{k}=\begin{bmatrix}I&0\end{bmatrix}$ then $Q_{k}^{\top}=S_{k}\mathcal{R}_{k}$. Thus, we can now write the following:
$v^{(i)}_{trg,t}$ $=$ $h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}\begin{bmatrix}I\\ 0\end{bmatrix}(-Q_{i}^{\top}h^{(i)}_{trg,t}+Q_{j}^{\top}h^{(j)}_{src,u})$ (29)
$v^{(i)}_{trg,t}$ $=$ $h^{(i)}_{trg,t}+\mathcal{R}_{i}^{\top}S_{i}^{\top}(-Q_{i}^{\top}h^{(i)}_{trg,t}+Q_{j}^{\top}h^{(j)}_{src,u})$ (30)
$v^{(i)}_{trg,t}$ $=$ $h^{(i)}_{trg,t}+Q_{i}(-Q_{i}^{\top}h^{(i)}_{trg,t}+Q_{j}^{\top}h^{(j)}_{src,u})$ (31)
And our proof is finished.

### v6

@@ B.2 Model Details
All artificial neural network models were implemented and trained using PyTorch (Paszke et al., 2019) on a single Nvidia Titan X GPU. Unless otherwise stated, all models used an embedding and hidden state size of 128 dimensions. To make the token predictions, each model used a two layer multi-layer perceptron (MLP) with GELU nonlinearities, with a hidden layer size of 4 times the hidden state dimensionality with 50% dropout on the hidden layer. The GRU and LSTM model variants each consisted of a single recurrent cell followed by the output MLP. Unless otherwise stated, the transformer architecture consisted of two layers using Rotary positional encodings (Su et al., 2023). Each model variant used the same learning rate scheduler, which consisted of the original transformer (Vaswani et al., 2017) scheduling of warmup followed by decay. We used 100 warmup steps, a maximum learning rate of 0.001 , a minimum of 1e-7, and a decay rate of 0.5. We used a batch size of 128, which caused each epoch to consist of 8 gradient update steps.
B.3 MAS (and Associated Variants) Training Details
@@
All MAS trainings were implemented and trained using PyTorch (Paszke et al., 2019) on single Nvidia Titan X GPUs. For each rotation matrix training, we use 10000 intervention samples and 1000 samples for validation and testing. We uniformly sampled corresponding indices upon which to perform interventions, excluding the B, T, and E tokens in the numeric equivalence tasks from possible intervention sample indices. In the Arithmetic task, we used the comma token for Rem Ops and Cumu Val interventions. When intervening upon a state in the demo phase in the numeric equivalence tasks, we uniformly sample a number of steps to continue the demo phase that will keep the object quantity below 20. We orthongonalize the matrices $Q_{\psi_{i}}$, using PyTorch’s orthogonal parametrization with default settings. PyTorch creates the orthogonal matrix as the exponential of a skew symmetric matrix. We train the rotation matrices for 1000 epochs, with a batch size of 512 used for each model index pairing. We only perform experiments considering two models. Each gradient step uses the average gradient over batches of all 4 $(i,j)$ ordered pairings. We select the checkpoint with the best validation performance for analysis. We use a learning rate of 0.001 and an Adam optimizer.
For the LSTM architecture, we perform MAS on a concatenation of the $h$ and $c$ recurrent state vectors (Hochreiter & Schmidhuber, 1997). In the GRUs, we operate on the recurrent hidden state. In the transformers, we operate on the residual stream following the first transformer layer (referred to as the Layer 1 Hidden States in Supplementary Figure 4).
Stepwise MAS applies Equation 3 to multiple, contiguous time-steps in the target and source sequences from step 0 to $u$. This allows for meaningful transfer of information in cases of anti-Markovian solutions (Grant et al., 2025). There is a question of what tokens to use to create each $h^{(trg)}_{\psi_{i}}$ before transfer. We show results where each target token comes from the original target sequence padded by the R token when $u$ exceeds the target sequence length. See Figure 1 for a visualization.
B.4 RSA Details
@@
We performed RSA on a subsample of a dataset of 15 sampled sequences for each object quantity ranging from 1-20 on the task that each model was trained on for each model. We first ran the models on their respective datasets to collect the latent representations. We sampled 1000 of these latent vectors as the sample representations in a matrix $M_{k}\in R^{N\times d_{k}}$ where $k$ refers to the model index, $N$ is the number of latent vectors ($N=1000$ in our analyses), $d_{k}$ is the dimensionality of a single latent vector for model $k$. We then calculated the sample cosine distance matrices (1-cosine similarity) for each model resulting in matrices $C_{k}\in R^{N\times N}$. Lastly we calculated the Spearman’s Rank Correlation Coefficient between the lower triangles of the matrices $C_{1}$ and $C_{2}$ as the RSA value using python’s SciPy package (Zar, 2005; Kriegeskorte et al., 2008; Virtanen et al., 2020). We resampled the 1000 vectors 10 times and recalculated the RSA score 10 times and report the average over these scores.
B.5 CKA Details
@@
(Kornblith et al., 2019). We resampled the 1000 vectors 10 times and recalculated the CKA score 10 times and report the average over these scores.
@@ B.6 Arithmetic
Arithmetic Task: This task consists of an indication of the number of addition/subtraction operations, an initial value, and then the operations interlaced with the cumulative value (cumu val) of the operations. An example sequence is: "B 3 4 + 3 = 7 , + 11 = 18 , - 5 = 13 E", where the first number indicates the number of operations in the trial. The 2nd number is a sampled start value. Each operand is a sampled value that is combined with the cumu val according to the operator. The number of operations is uniformly sampled from 1-10. The starting value is uniformly sampled from 0-20. All numeric values are restricted inclusively to 0-20. The operations are uniformly sampled from $\{+,-\}$ when the cumu val is in the range 1-19. When 0 and 20 the operations are "+" and "-" respectively. The operands are uniformly sampled from the set that restricts the next cumu val to the range 0-20. The cumu val is shown after the "=" for each operation. The "," token follows the cumu val until the end of the sequence in which the E token replaces the comma. The model must predict the "=" tokens, cumu vals, commas, and the E token for a trial to be correct. We use a base 21 token system so that all values correspond to a single token.